If you note the following yield curve in The Wall Street Journal, what is the one-year forward rate for the period beginning one year from today, 2f1 according to the unbiased expectations theory? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))

Maturity Yield
One day 2.00 %
One year 5.50
Two years 6.50
Three years 9.00
Answer #1

Two year forward rate = (1+ Two year bond’s interest rate)^2/ (1+ One years interest rate) -1

= (1+0.065)^2/ (1+0.055) -1

=1.075095 -1

= 7.51%